OpenStudy (anonymous):

Let g(x)=2x and h(x)=x^2+4. Find the value (g o h) (0)

OpenStudy (anonymous):

so (g o h)(0) means g(h(0))

OpenStudy (anonymous):

first figure out h(0), than plug that into g for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Like this? 0=0^2+4

OpenStudy (anonymous):

-4=0??

OpenStudy (anonymous):

I think you mean h(0). h(0) = 0^2 + 4 h(0) = 0 + 4 h(0) = 4

OpenStudy (anonymous):

Ah ok

OpenStudy (anonymous):

so now we know h(0) = 4 so we can do g(h(0)) = g(4) right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

g(4) = 2(4) g(4) = 8

OpenStudy (anonymous):

so (g o h)(0) = 8

OpenStudy (anonymous):

Ok great much easier to understand thank you

OpenStudy (anonymous):

you're welcome